2. A classic pair of complementary operators are £ and Pz. In class we showed these operators obeyed the Heisenberg relationship for the harmonic oscillator. Show they also obey this relationship for the particle in the box. Note: your answer will be a fiunction of the quantum number n. Show the relationship holds for n - I and any larger value ofn.
Answer #1

Energy of a particle in a box (1D infinite potential) is:

E_n = rac{n^2h^2}{8mL^2} = rac{n^2hbar^2pi^2}{2mL^2}

but En-2:n

so, rac{p_n^2}{2m} = rac{n^2hbar^2pi^2}{2mL^2}

=> p_n^2 = rac{n^2hbar^2pi^2}{L^2}

=> 71

since nhT 2 for every n [n = 1,2,3,4,.....] and 0 < x < L

therefore, Delta x Delta p leq rac{hbar}{2} is satisfied for every n.